# GLCF: A Global-Local Multimodal Coherence Analysis Framework for Talking Face Generation Detection

### v1

@@ Introduction
Figure 2: The construction of 11 scenarios. The green box indicate inputs from the same source, while the red box is the opposite. Green audio denotes genuine audio, whereas red audio signifies forgery audio. Blank means no input.
@@
Figure 3: The distribution of training set.
Figure 4: The distribution of testing set.
@@
In figure 4 and figure 4, we present the generation distribution of the dataset on each generation method. To simulate the detection of potential novel real-world scenarios, we add a small portion of the LSR2 dataset to our testing set, which remains unexposed to the training phase, thereby increasing the complexity of the presented scenes. Given the rapid advancements in generative artificial intelligence and talking face generation technologies, we incorporate the IP_ LAP, which also remains unexposed to the training set, to generate a small portion of testing cases. This cross-domain evaluation emphasizes the necessity for detection methods to exhibit generalization abilities to ensure their superior performance within the testing set. Hence, these configurations facilitate the exploration of generalized, adaptable detection methods, enabling us to effectively tackle unknown and sophisticated forgery challenges.
Furthermore, we combine videos of different shooting scenes and audio of different topics to simulate more semantic scenarios. The detailed distribution is shown in Figure 6 and Figure 6. Totally about 40 kinds of scenarios.
@@
Figure 5: The detailed setting of the SadTalker method.
Figure 6: The detailed setting of the other three methods. R and F respectively stand for real and fake audio, and the number stands for the input consistency levels, eg: R1 – audio and reference are same. The legend is the audio set. The center is image and video datasets.
@@
Audio Stream: We use a pre-trained wav2vec model and a two-layer residual network to map the input audio stream into the feature space, obtaining $F_{a}$. The output will be fed into two streams. One of them introduces a complete forgery audio detection stream, which directly predicts the authenticity of the audio. This stream also facilitates the addition of the audio loss function $L_{a}$, better fine-tuning of Wav2Vec and training of ResLayer. The total loss function is:
@@
Video-Audio Fusion Module: We design an audiovisual fusion module based on a multi-head cross-attention mechanism, the main purpose of which is to detect the coherence between audiovisual modes. Specifically, We input audio features into a 1D CNN to align the feature dimension with the visual feature vector, and then further input a 3D CNN to obtain $Query\in R^{C\times\ T\times H\times W}$. Similarly, visual features are entered into the 3D CNN to get the $Key\in R^{C\times\ T\times H\times W}$. $Value$ is a linear combination of audiovisual features, and the entire calculation process can be expressed as:
@@
Datasets: To evaluate our dataset and methodology, we conduct experiments on our MSTF and three other challenging datasets: FaceForensis++, FakeAVCeleb, and DFDC.
@@
Evaluation metrics: In our experiments, we use the accuracy rate (ACC) for evaluation.
Implementation Details: We use MTCNN to detect faces in all datasets and then crop them. We adjust the face images to 256*256 to retain more facial information. We use NVIDIA A100 Tensor Core GPU.
@@
Intra-Dataset Comparisons.
@@
Dataset: Experimental results show that our raw MSTF dataset is even harder to detect content from than the well-known and challenging FF++ dataset(C40), revealing the necessity of specialized detection for talking face generation and the effectiveness and validity of the MSTF dataset in facilitating such endeavors.
@@
Method: Our method achieve the best performance on the FakeAVCeleb, DFDC, and MSTF datasets. On the MSTF dataset, it surpasses the second-best performer, MiNet, by 2.37%. For the FakeAVCeleb dataset, focusing on the binary classification of genuine and fake samples, our method also outperforms the second-best performer, Multi-att, by 0.59%. Similarly, on DFDC, we outperform F3Net, the second-best performer, by a narrow margin of 0.13%. Experimental results demonstrate that our methodology is not merely confined to the specific domain of talking face generation detection but also exhibits a significant degree of generalizability, enabling it to adeptly address a wide array of forgery tactics, including traditional deepfake.
@@ Acknowledgments
Appendix
Image and Video Collection
For video data, we randomly extract one frame and save it as image data. CelebA-HQ: The dataset consists of 30,000 images with 1024*1024 resolution.We randomly selected 20,000 images from it.Each one of them is high-quality and with the face centered.
Cream-D: We have selected 10,659 images of frontal faces as the reference input for emotion recognition.
Vox: VoxCeleb was filmed in a variety of settings, we selected 18,000 videos with better quality and more centered speaker faces from it, and used a pre-trained face recognition model to perform gender classification and face cropping to adapt to the needs of different generative models and achieve better generation results.
Vox2: VoxCeleb2 contains one million speech videos of more than 6,000 celebrities. For each celebrity, we selected two videos, for a total of about 12,000 videos, and face cropped the videos. Meanwhile, we randomly selected 300 videos of celebrities for generating the test set, which will not appear in the training set.
Mead: We selected 18 actors’ videos for generating the training set and 4 for generating the test set, which guarantees the character independence of the test and training sets. To ensure the quality of the generation, we finally selected only 16,000 videos containing the front face angle, of which 12,000 were used for the training set and 4,000 for the test set.
DFDC: We use 4464 separate training videos and 780 test videos from DFDC, each of which is 15 seconds in length. We use the real videos only.
LSR2: In order to enrich the scenarios of the test data, we selected 1000 videos from the lip-reading dataset LSR2, and these were only used for the generation of the test data.
Youtobe video: We also collected speech videos from youtobe with some high quality videos, including scenes from news broadcasts and interviews.
Audio Collection
Audio of FakeAVceleb: We extract all the forged audio from this dataset to increase the variety of forged audio.
LibriSpeech: We selected more than 2,400 audio clips for each gender, each with a length greater than 5 seconds and less than 15 seconds.
ASVspoof2019: A total of 4,000 forged audio samples were selected from the evaluation set for inclusion in the training set, while 800 samples were chosen for the test set. The selected samples were balanced in terms of gender and had a duration of at least five seconds, representing 13 distinct forms of forgery.
FakeVox: We respectively generate an emotional and a standard audio datasets, each containing 2000 samples.
Figure 11: The detailed of the spatial aggregation of DCTAM.
Spatial Aggregation
As shown in Figure 11, We supplement the detailed design diagram of spatial aggregation in DCTAM.

### v2

@@ Introduction
Figure 2: The construction of 11 scenarios. The six boxes in the above big image correspond to the six boxes in each small image below. The green box indicates inputs from the same source, while the red box is the opposite. Green audio denotes genuine audio, whereas red audio signifies forgery audio. Blank means no input.
@@
In figure 3, we present the generation distribution of the dataset on each generation method. To simulate the detection of potential novel real-world scenarios, we add a small portion of the LSR2 dataset to our testing set, which remains unexposed to the training phase, thereby increasing the complexity of the presented scenes. Given the rapid advancements in generative artificial intelligence and talking face generation technologies, we incorporate the IP_ LAP, which also remains unexposed to the training set, to generate a small portion of testing cases. This cross-domain evaluation emphasizes the necessity for detection methods to exhibit generalization abilities to ensure their superior performance within the testing set. Hence, these configurations facilitate the exploration of generalized, adaptable detection methods, enabling us to effectively tackle unknown and sophisticated forgery challenges.
Furthermore, we combine videos from different shooting scenes and audios on different topics to simulate more than 20 semantic scenarios, resulting in a total of about 40 different scenarios when combined with generation scenarios.
@@
Audio Stream
We use a pre-trained wav2vec model and a two-layer residual network to map the input audio stream into the feature space, obtaining $F_{a}$. The output will be fed into two streams. One of them introduces a complete forgery audio detection stream, which directly predicts the authenticity of the audio. This stream also facilitates the addition of the audio loss function $L_{a}$, better fine-tuning of Wav2Vec and training of ResLayer. The total loss function is:
@@
Video-Audio Fusion Module
We design an audiovisual fusion module based on a multi-head cross-attention mechanism, the main purpose of which is to detect the coherence between audiovisual modes. Specifically, We input audio features into a 1D CNN to align the feature dimension with the visual feature vector, and then further input a 3D CNN to obtain $Query\in R^{C\times\ T\times H\times W}$. Similarly, visual features are entered into the 3D CNN to get the $Key\in R^{C\times\ T\times H\times W}$. $Value$ is a linear combination of audiovisual features, and the entire calculation process can be expressed as:
@@
Datasets
To evaluate our dataset and methodology, we conduct experiments on our MSTF and three other challenging datasets: FaceForensis++, FakeAVCeleb, and DFDC.
@@
Evaluation metrics
In our experiments, we use the accuracy rate (ACC) for evaluation.
Implementation Details
We use MTCNN to detect faces in all datasets and then crop them. We adjust the face images to 256*256 to retain more facial information. We use NVIDIA A100 Tensor Core GPU.
@@
Intra-Dataset Comparisons
@@
Experimental results show that our raw MSTF dataset is even harder to detect content from than the well-known and challenging FF++ dataset(C40), revealing the necessity of specialized detection for talking face generation and the effectiveness and validity of the MSTF dataset in facilitating such endeavors.
@@
In the meantime, our method achieves the best performance on the FakeAVCeleb, DFDC, and MSTF datasets. On the MSTF dataset, it surpasses the second-best performer, MiNet, by 2.37%. For the FakeAVCeleb dataset, focusing on the binary classification of genuine and fake samples, our method also outperforms the second-best performer, Multi-att, by 0.59%. Similarly, on DFDC, we outperform F3Net, the second-best performer, by a narrow margin of 0.13%. Experimental results demonstrate that our methodology is not merely confined to the specific domain of talking face generation detection but also exhibits a significant degree of generalizability, enabling it to adeptly address a wide array of forgery tactics, including traditional deepfake.